OpenStudy (anonymous):

I've tried solving this and everytime I get the wrong answer. can someone help me? What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 0 4 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system.

OpenStudy (anonymous):

use substitution or elimination.

OpenStudy (anonymous):

I did and got the wrong answer

OpenStudy (anonymous):

I got it all the way to 6x=-2y+36

OpenStudy (anonymous):

@Jhannybean can you help me?

OpenStudy (anonymous):

seems like it is infinite solutions.

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

2x − y = 6 4x − 2y = 12 2 -1 6 4 -2 12 2 -1 6 2 -1 6 4 -2 12 -2r1+r2 0 0 0 solve 2 - 6 for y.

OpenStudy (anonymous):

can you help me with this one? What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 −2 −3 x can be any number as there are infinitely many solutions to the system. There is no x value as there is no solution to the system --------------------------------------------------------------------------------

OpenStudy (anonymous):

lets first do the augmented matric just get the coefficients and right hand side. 3 -6 3 7 -5 -11 with me so far?

OpenStudy (anonymous):

yes, then what?

OpenStudy (anonymous):

we try elimination. Lets try to get a 0 where the 7 is at. 3 -6 3 3 -6 3 7 -5 -11 -7/3 r1+r2 0 9 -18 If i did my calculations right. WIth me so far?

OpenStudy (anonymous):

alright, yeah

OpenStudy (anonymous):

3 -6 3 3 -6 3 7 -5 -11 -7/3 r1+r2 0 9 -18 now we try to make the 9 a 1 so r2*1/9 = 3 -6 3 0 1 2 as you can see that in row 2 if we go back to an equation. y=2 :D

OpenStudy (anonymous):

so its -2?

OpenStudy (anonymous):

2 :P silly

OpenStudy (anonymous):

-2 :P sorry

OpenStudy (anonymous):

forgot that it was -18 not just 18 :P good catch smarty :D

OpenStudy (anonymous):

thank you :) what about another one? :)

OpenStudy (anonymous):

sure :D

OpenStudy (anonymous):

oh btw it has one unique solution :D

OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x + y = 14 –3x + 2y = 1 5 3 y can be any number as there are infinitely many solutions to the system. There is no y value as there is no solution to the system.

OpenStudy (anonymous):

do the augmented matrix for me first :) see if you know that.

OpenStudy (anonymous):

I have no idea what that is

OpenStudy (phi):

I think they want you to add the 2 equations

OpenStudy (anonymous):

getting rid of the variables. Just focus on numbers

OpenStudy (anonymous):

3x + y = 14 –3x + 2y = 1 is 3 1 14 -3 2 1

OpenStudy (anonymous):

yeah they want me to solve for the first equation then plug that one into the second one and solve for that

OpenStudy (phi):

try adding the 2 equations.

OpenStudy (anonymous):

its 5, that was easy lol

OpenStudy (anonymous):

yup lol

OpenStudy (anonymous):

thank you!